OpenStudy (anonymous):

Determine the best method to solve each system of equations. How would I solve this? 5x + 3y = 16 3x - 5y = -4 Thanks!

OpenStudy (anonymous):

elimination right? or substitution?

OpenStudy (anonymous):

Which ever works is fine

OpenStudy (anonymous):

Substitution is the only one i can think of. 15x + 9y = 16 15x = 16 - 9y 15x - 25y = -4 15x = 25y - 4 25y - 4 = 16 - 9y 34y - 4 = 16 34y = 20 y = 20/34 = 10/17 15x = 16 - (10/17) * 9x 15x = 16 - (90/17) * x 15x + (90/17) * x = 16 x((255/17) + (90/17)) = 16 x(335/17) = 16 335x = 112 x approximately equals 3, or 2.9910714285714285714285714285714

OpenStudy (anonymous):

Thanks for the help!